OpenStudy (anonymous):

Solve the system of equations. x + y + z = 18 3x + 4y + 5z = 72 x + z = 8

OpenStudy (anonymous):

hmmm no clue :3 srry mate

OpenStudy (anonymous):

is there something missing in the last one or is it just \(x+z=8\)?

OpenStudy (anonymous):

@satellite73 - it's just x + z = 8

OpenStudy (anonymous):

yeah it is probably what you wrote this is donkey work, let the computer do it http://www.wolframalpha.com/input/?i=+x+%2B+y+%2B+z+%3D+18%2C+3x+%2B+4y+%2B+5z+%3D+72%2C+x+%2B+z+%3D+8

OpenStudy (anonymous):

looks like you get \((4,10,4)\)

OpenStudy (anonymous):

we can do it by hand if you like, but it is not worth the aggravation imho

OpenStudy (anonymous):

Thanks, but to get credit for the problem in my class, I need to show my work. Sorry.

OpenStudy (anonymous):

ok then multiply the first equation by \(-4\) and add it to the second

OpenStudy (anonymous):

\[ x + y + z = 18\\ 3x + 4y + 5z = 72\] becomes \[-4x-4y-4z=-72\\ 3x+4y+5z=72\]

OpenStudy (anonymous):

add them up, get \[-x+z=0\] and you also have the last equation \[x+z=8\] solve \[-x+z=0\\ x+z=8\] by addition, get \[2z=8\\ z=4\]

OpenStudy (anonymous):

now that you have \(z\) you can find \(x\) easily, then you can find \(y\)

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

@satellite73 - I don't think so.... Sorry. Let's say I plug in z into the first equation. I would get x + y = 14. What do I do with this equation?

OpenStudy (anonymous):

we're good with \(z=4\) right?

OpenStudy (anonymous):

the last equation is \(x+z=8\) so if \(z=4\) we have \[x+4=8\\x=4\]

OpenStudy (anonymous):

now that you have \(x=4,z=4\) put those in either of the first two equations to find \(y\)

OpenStudy (anonymous):

@satellite73 - Oh okay, I get it now. If I use the first equation; x + y + z = 18 and plug in x and z, 4 + y + z = 18 y = 10.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

@satellite73 - Thank you so much!

OpenStudy (anonymous):

yw